Time table for SS 08

09:00 - 10:00

10:00 - 11:00

11:00 - 12:00

12:00 - 13:00

13:00 - 14:00

14:00 - 15:00

15:00 - 16:00

16:00 - 17:00

17:00 - 18:00 Monday Tuesday Wednesday Thursday

Friday

Differential Equations in Image Processing and Computer Vision

Weickert

E13 001

Bachelor Thesis Bachelor Thesis

Processing of Matrix Valued Images

Didas E11 306

Differential Equations in Image Processing and Computer Vision Tutorial

Mainberger {E13 \atop SR16/104}

Differential Equations in Image Processing and Computer Vision

Weickert

E13 001

Bachelor Thesis

Oberseminar MIA

Weickert E11 306